# LLM4Drive: A Survey of Large Language Models for Autonomous Driving

### v3

@@ 4 Datasets in LLM4AD
Traditional datasets such as nuScenes dataset Caesar et al. (2019); Fong et al. (2021) lack action description, detailed caption, and question-answering pairs which are used to interact with LLMs. The BDD-x Kim et al. (2018), Rank2Tell Sachdeva et al. (2023), DriveLM Sima et al. (2023), DRAMA Malla et al. (2023), NuPrompt Wu et al. (2023) and NuScenes-QA Qian et al. (2023) datasets represent key developments in LLM4AD research, each bringing unique contributions to understanding agent behaviors and urban traffic dynamics through extensive, diverse, and situation-rich annotations. We give a summary of each dataset in Table 1. We give detailed descriptions below.

### v4

@@ 4 Datasets in LLM4AD
Traditional datasets such as nuScenes dataset Caesar et al. (2019); Fong et al. (2021) lack action description Lu et al. (2024), detailed caption, and question-answering pairs which are used to interact with LLMs. The BDD-x Kim et al. (2018), Rank2Tell Sachdeva et al. (2023), DriveLM Sima et al. (2023), DRAMA Malla et al. (2023), NuPrompt Wu et al. (2023) and NuScenes-QA Qian et al. (2023) datasets represent key developments in LLM4AD research, each bringing unique contributions to understanding agent behaviors and urban traffic dynamics through extensive, diverse, and situation-rich annotations. We give a summary of each dataset in Table 1. We give detailed descriptions below.
@@
NuInstruct Ding et al. (2024): It is a dataset featuring 91K multi-view video-QA pairs spanning 17 subtasks, each requiring comprehensive information such as temporal, multi-view, and spatial data, thereby significantly raising the complexity of the challenges. It developed a SQL-based method that automatically generates instruction-response pairs, inspired by the logical progression of human decision-making.
OpenDV-2K Yang et al. (2024): This dataset is a large-scale multimodal dataset for autonomous driving, comprising 2059 hours of curated driving videos, including 1747 hours from YouTube and 312 hours from public datasets, with automatically generated language annotations to support generalized video prediction model training.